Federal Reserve Bank of Minneapolis Research Department Staff Report 250

Revised August 1998

Explaining Cross-Country Income Differences^{†}

Ellen R. McGrattan Federal Reserve Bank of Minneapolis

James A. Schmitz, Jr. Federal Reserve Bank of Minneapolis

ABSTRACT This chapter reviews the literature that tries to explain the disparity and variation of GDP per worker and GDP per capita across countries and across time. There are many potential explanations for the different patterns of development across countries, including differences in luck, raw materials, geography, preferences, and economic policies. We focus on differences in economic policies and ask to what extent can differences in policies across countries account for the observed variability in income levels and their growth rates. We review estimates for a wide range of policy variables. In many cases, the magnitude of the estimates is under debate. Estimates found by running cross-sectional growth regressions are sensitive to which variables are included as explanatory variables. Estimates found using quantitative theory depend in critical ways on values of parameters and measures of factor inputs for which there is little consensus. In this chapter, we review the ongoing debates of the literature and the progress that has been made thus far.

† The views expressed herein are those of the authors and not necessarily those of the Federal Reserve Bank of Minneapolis or the Federal Reserve System. We thank Francesco Caselli, Bill Easterly, Boyan Jovanovic, Pete Klenow, Narayana Kocherlakota, and Ed Prescott for helpful comments. Data used in this chapter are available at our web site: http://research.mpls.frb.fed.us/research/sr.